seema is 4 years younger than her brother Shiv.After 5 years their ages will be in the ratio of 3 is to 4.Find their present ages.
Answer by Maths68(1474)

Let
Present age of Seema = s years old
Present age of Shive = h years old


Seema is 4 years younger than her brother Shiv
s=h-4.............(1)

After 5 years
Seema will be = s+5 years old
Shive will be = h+5 years old
Then
Their ages will be in the ratio of 3 is to 4
(s+5)/(h+5)=3/4
4(s+5)=3(h+5)
4s+20=3h+15
4s-3h=15-20
4s-3h=-5...................(2)
Substitute the value of s from (1) to (2)
4(h-4)-3h=-5
4h-16-3h=-5
4h-3h=-5+16
h=11
Now substitute the value of h in (1)
s=h-4.............(1)
s=11-4
s=7


Present age of Seema = s = 7 years old
Present age of Shive = h = 11 years old